# An Inverse Problem for the Prescribed Mean Curvature

### v1

@@ Abstract.
We extend the study of inverse problems for minimal surfaces by considering the inverse source problem for the prescribed mean curvature equation
@@
We prove that in two dimensions, the source function $H$ is uniquely determined by the associated Dirichlet-to-Neumann map. A notable feature of this problem is that although the equation is posed on an Euclidean domain, its linearization yields an anisotropic conductivity equation where the coefficient matrix corresponds to a Riemannian metric $g$ depending on the background solution. This work represents the first treatments of inverse source problems for quasilinear equations.
The proof relies on the higher order linearization method. The main methodological contribution is the development of an approach to decouple the resulting system of nonlinear algebraic and geometric equations, which enables the complete determination of the source term from boundary measurements. The decoupling uses a Liouville type uniqueness result for conformal mappings.
@@ 1.2. Earlier works
Our work is a new addition to a growing branch of inverse problems for nonlinear partial differential equations. In [Isa93] the author linearized the DN map to solve an inverse problem for a nonlinear PDE. By linearizing one can then use methods for linear equations to solve the problem. In recent years the interest for these has grown and it started with the works [KLU18, FO20a, LLLS21a]. In these works the method of higher order linearization was introduced. In this method one linearizes the DN map multiple times and solves an inverse problem at each step, gaining new information about the original problem.
@@
Inverse problems for elliptic second order semilinear equations have been considered in [IS94, IN95, Sun10, IY13, LLLS21a, LLLS21b, LLST22, KU20, FO20b, FLL23, ST23, Nur23a, JNS23, JNS25] and inverse problems for nonlinear magnetic Schrödinger equations in [KU23, LZ23, Ma20, KUY24]. We also mention the works for nonlinear biharmonic equations [BKSU25, NS25], $p$-harmonic equation [BKS15, BIK18, CF25] and the works for nonlinear fractional equations [Lin22, LL19a, LL23, LL19b]. See also [Las25] for a recent survey on this subject.
@@ 4.2. Showing that $\phi$ is the identity
That is $\phi$ is a conformal mapping in Euclidean space. Combining this with $\phi|_{\partial\Omega}=\mathrm{Id}$ a Liouville typer result (see e.g. [Lio97, Proposition $3.$]) gives that $\phi$ is the identity in $\Omega$.
@@ 6. Proving that $T=0$
Acknowledgment. The authors would like to thank Niko Jokela for helpful discussion. T.L. was supported by the Academy of Finland (Centre of Excellence in Inverse Modeling and Imaging, grant numbers 284715 and 309963). J.N. was supported by the Finnish Centre of Excellence in Inverse Modelling and Imaging (Academy of Finland Grant 284715), the Research Council of Finland (Flagship of Advanced Mathematics for Sensing Imaging and Modelling grants 359208 and 359183) and by the Emil Aaltonen Foundation.

### v2

@@ Abstract.
We extend the recent study of inverse problems for minimal surfaces by considering the inverse source problem for the prescribed mean curvature equation
@@
This work also represents the first treatment of inverse source problems for quasilinear equations. We prove that in two dimensions, the source function $H$ is uniquely determined by the associated Dirichlet-to-Neumann map. A notable feature of this problem is that although the equation is posed on an Euclidean domain, its linearization yields an anisotropic conductivity equation where the coefficient matrix corresponds to a Riemannian metric $g$ depending on the background solution.
The main methodological contribution is the derivation of a coupled nonlinear system of algebraic and geometric partial differential equations from boundary measurements. Similar systems will naturally appear in other inverse problems for quasilinear equations. We solve the system using a Liouville type uniqueness result for conformal mappings, which recovers the source function uniquely.
@@ 1.2. Earlier works
Our work is a new addition to a growing branch of inverse problems for nonlinear partial differential equations. In [Isa93] the author linearized the DN map to solve an inverse problem for a nonlinear PDE. By linearizing one can then use methods for linear equations to solve the problem. In recent years the interest for these has grown and it started with the works [KLU18, FO20, LLLS21a]. In these works the method of higher order linearization was introduced. In this method one linearizes the DN map multiple times and solves an inverse problem at each step, gaining new information about the original problem.
@@
Inverse problems for elliptic second order semilinear equations have been considered in [IS94, IN95, Sun10, IY13, LLLS21a, LLLS21b, LLST22, KU20, FO20, FLL23, ST23, Nur23a, JNS23, JNS25] and inverse problems for nonlinear magnetic Schrödinger equations in [KU23, LZ23, Ma20, KUY24]. We also mention the works for nonlinear biharmonic equations [BKSU25, NS25], $p$-harmonic equation [BKS15, BIK18, CF25] and the works for nonlinear fractional equations [Lin22, LL19a, LL23, LL19b]. See also [Las25] for a recent survey on this subject.
@@ 2. First linearization
Remark 2.3.
We note that prescribed Jacobian equations such as (2.14) may lack unique solutions [CDK11, Section 1.4]. To resolve this ambiguity, we proceed to the second linearization in Section 4, which yields additional constraints on $\phi$. Ultimately, these constraints force $\phi$ to be a conformal mapping, whose uniqueness then follows from Liouville-type rigidity.
@@ 4.2. Showing that $\phi$ is the identity
That is $\phi$ is a conformal mapping in Euclidean space. Combining this with $\phi|_{\partial\Omega}=\mathrm{Id}$ a Liouville type result (see e.g. [Lio97, Proposition $3.$]) gives that $\phi$ is the identity in $\Omega$.
@@ 6. Proving that $T=0$
Acknowledgment. The authors would like to thank Niko Jokela and Gyula Csató for helpful discussion. T.L. was supported by the Academy of Finland (Centre of Excellence in Inverse Modeling and Imaging, grant numbers 284715 and 309963). J.N. was supported by the Finnish Centre of Excellence in Inverse Modelling and Imaging (Academy of Finland Grant 284715), the Research Council of Finland (Flagship of Advanced Mathematics for Sensing Imaging and Modelling grants 359208 and 359183) and by the Emil Aaltonen Foundation.